OpenStudy (anonymous):

During a 4th of July weekend, 32 vehicles became trapped on the Sunshine Skyway Bridge while it was being repaved. A recent city ordinance decreed that only cars with 4 wheels and trucks with six wheels could be on the bridge at any given time. If there were 148 tires that needed to be replaced to due to damage, how many cars and trucks were involved in the incident? i got this far: c = number of cars; t = number of trucks c + t = 32 4c + 6t = 148

OpenStudy (anonymous):

i solved for t and got 10 but when i solve for c i get 14 and that doesnt add upto 32 :(

OpenStudy (anonymous):

Could you please shoe me the steps you did?

OpenStudy (anonymous):

if you found dat c=10, den from first equation.......... t must be 22, where did you get 14 from?

OpenStudy (anonymous):

wait sorry, you cant get t as 10, show us ur working

OpenStudy (anonymous):

for t... 4(32-t) +6t = 148 128 -4t+6t = 148 2t = 20 t =10 then for c... 4(10) + 6(32-c) = 148 40 + 192 - 6c = 148 232 -6c = 148 -6c = -84 6c = 84 c =14

OpenStudy (anonymous):

ive checked over it a bunch of times and its not working

OpenStudy (anonymous):

after you get t substitute it into first equation

OpenStudy (anonymous):

thats wat i did @Tushara

OpenStudy (anonymous):

t=10 then all you do is using c+t=32 c+10=32 c=22 dont do the long process

OpenStudy (anonymous):

t is correct! Well done! Now we can use the 1st equation instead of the 2nd: c + t = 32 c+10=32 So, what do we get for c?

OpenStudy (anonymous):

@rafabc02....you'd get 22 for c but why doesnt it work in the second equation??

OpenStudy (anonymous):

what you did was : 4(10) it is 4c so you cant substitute the t value into c c is not 10

OpenStudy (anonymous):

OHHHH

OpenStudy (anonymous):

It does work. Let's see: t=10, c=22. So: 4c + 6t = 148 4*22+6*10=88+60=148 :)

OpenStudy (anonymous):

thanks soo much @Tushara

OpenStudy (anonymous):

yes :) np

OpenStudy (anonymous):

wow.....tht was an epic fail...i feel stupid now :P

OpenStudy (anonymous):

haha silly mistakes happen

OpenStudy (anonymous):

haha yea

OpenStudy (anonymous):

You don't have to feel stupid! Everyone makes mistakes! No worries!

OpenStudy (anonymous):

alright guys...thanks so much for the help....makin my life easier :D .... bye! :)

OpenStudy (anonymous):

You are welcome! Have a wonderful day!

OpenStudy (anonymous):

aite u too!

OpenStudy (anonymous):

Thanks!